OpenStudy (anonymous):

Find f(-3) if f(x) = x 2. -9 -6 6 9

OpenStudy (anonymous):

b

Nnesha (nnesha):

f(-3) means x=-3 so substitute -3 for x

Nnesha (nnesha):

what did you get ?

Nnesha (nnesha):

nice don't answer it you will get the result !! :=) good luck!

OpenStudy (unklerhaukus):

\[f(x) = x^2\\ f(-3)= (-3)^2= -3\times-3 =\]